# Large Language Models for Design Structure Matrix Optimization

### v1

@@ 5.2 Convergence Analysis with Stochastic Algorithms
We evaluated the convergence behavior of our LLM-based methods in comparison with three configurations of GA, as shown in Figure 3. Due to their differing operational mechanisms, each GA iteration explores multiple unique solutions based on population dynamics and stochastic operations. In contrast, our LLM-based approach generates one unique solution per iteration. To ensure a fair comparison, we standardized convergence speed in terms of the number of unique solutions explored, rather than iteration count. Although the maximum generation number for GA was set to 2,000 (refer to Appendix 2), the comparison in Figure 3 focuses on the first 10,000 unique solutions to improve clarity of visualization.
Across all four DSM cases, our methods consistently demonstrate superior convergence speed relative to all GA variants. The LLM-based methods reach significantly lower objective values with far fewer explored sequences, highlighting their efficiency. This convergence advantage is evident across all cases. In each scenario, our method identifies a relatively high-quality initial solution and then continues to refine it through iterative improvement. This supports the hypothesis that LLMs can perform efficient in-context learning, adapting and optimizing based on previously observed sequences.
Moreover, the results further highlight the value of contextual domain knowledge in improving the convergence behavior. In Figures 3A (UCAV), 3C (Heat Exchanger), and 3D (Brake System), the version of our method that incorporates domain knowledge (red line) consistently outperforms the variant without such knowledge (gray line), both in terms of convergence speed and final solution quality. This indicates that LLMs are capable of effectively combining semantic reasoning with structural optimization to solve CO problems more intelligently. An interesting exception appears in Figure 3B (Microfilm Cartridge), where both variants of the LLM-based method reach the optimal solution in the first step. As a result, the two lines overlap entirely, reflecting identical performance in this case. This suggests that, for some DSMs with relatively less complexity or clearer structure, even minimal attempts may suffice for the LLM to infer optimal solutions.
@@ 5.4 Implications for Engineering Design
Figure 4: Optimization trajectory on the Heat Exchanger DSM case using the proposed method
@@
In sum, the proposed method can directly contribute to design planning and system-level coordination in engineering design practice. For example, in activity-based DSMs, it supports early-phase decision-making by suggesting task sequences that mitigate rework and streamline communication. In parameter-based DSMs, it reduces technical complexity by identifying sequencing patterns that minimize feedback loops. By leveraging both system structure information and domain knowledge, our method enables the LLM to reason in alignment with engineering logic, rather than relying solely on statistical correlations. This results in more effective and efficient optimization outcomes, ultimately accelerating and enhancing the overall engineering design process.
While this study focuses on sequencing problems within DSMs, the proposed approach can be generalized to a broader class of CO problems [57, 58, 59] in engineering and more fields. A more generalized framework is illustrated in Figure 5. During the engineering design process, systems or tasks can often be formulated as iterative optimization problems over formal representations, where multimodal information (such as task descriptions, parameter interactions, or architectural constraints encoded in text, diagrams, or code format) is encoded and integrated. The framework captures this pattern by transforming such information into a combined input information and integrating them with objective evaluation in a closed-loop reasoning cycle, enabling generation-based optimization using foundational AI models.
Figure 5: Foundational AI Model-based Framework for Solving General CO Problems
@@ 6 Limitations and Future Work
Although the proposed method exhibits strong performance, several limitations remain. First, the current evaluation is based on a relatively small number of DSM cases. Expanding the dataset to include larger and more complex systems from diverse engineering domains would allow for more comprehensive analysis and improved statistical reliability. Second, this study focuses specifically on sequencing problems, whereas the proposed framework may be applicable to a wider range of CO problems in engineering design, such as modularization and change propagation analysis [57, 58, 59]. As presented in Figure 5, the framework is designed to generalize beyond sequencing by incorporating structured system information and domain-specific semantics into an iterative optimization process. Future work may explore the effectiveness and adaptability of this approach across these expanded tasks.
Another important direction for future research lies in improving the interpretability of the optimization process. In this paper, we presented a preliminary case illustrating how the LLM refines the DSM through a sequence of iterative updates. However, the underlying decision-making and reasoning mechanisms of LLMs still remain relatively opaque. Researchers may explore more effective techniques for LLM interpretability, such as integrating complementary models, enabling self-explanation, and visualizing decision pathways [60].
Additionally, the current framework primarily relies on textual representations of domain knowledge and system structures. While effective, this may limit its ability to fully leverage non-textual modalities such as diagrams, tables, or structured data formats commonly used in engineering practice. In future work, we will explore Multimodal Large Language Models (MLLMs) [61] to incorporate visual or tabular information, enriching the reasoning input and expanding the framework’s applicability to more complex engineering workflows.
@@ 7 Conclusion
In this paper, we have proposed a knowledge-informed optimization approach that leverages LLMs to perform iterative combinatorial optimization on DSMs. By integrating structural network information with contextual domain knowledge, the proposed method leverages the advanced reasoning capabilities of LLMs to iteratively generate and refine high-quality solutions. Experiments conducted on various DSM cases demonstrate that the proposed approach outperforms both stochastic and deterministic baselines in terms of convergence speed and solution quality. The consistent performance gains observed when incorporating domain knowledge further reveal the value of combining semantic insight with mathematical reasoning. These results suggest that LLMs provide a promising foundation for a new paradigm in design optimization and decision support, particularly within the complex and knowledge-intensive environments of engineering design.
@@ Appendix 1. Full prompts
Prompt for the proposed approach (removing knowledge):
@@
Appendix 2. Parameter settings for three variants of GA
In all three variants of the GA used for DSM sequencing tasks, we employed the following shared settings: the number of generations was set to 2,000, with a selection mechanism using tournament selection and mutation using shuffled indexes. The GA was implemented using the DEAP library with standard configurations for initial population generation. The different configurations for each GA setting are shown in the table below. The meaning of each parameter also refers to [49].

### v2

@@ 5.2 Convergence Analysis with Stochastic Algorithms
We evaluated the convergence of our LLM-based methods in comparison with three variants of GA and the OmeGA algorithm, as shown in Figure 4. Because of their differing operational mechanisms, each GA iteration evaluates a population of candidate solutions generated through stochastic variation operators. Specifically, each GA generation explores 50 unique solutions, corresponding to its population size. Similarly, each generation of the OmeGA algorithm evaluates 50 unique solutions per block, followed by additional evaluations during the local search phase for the best individual in each generation. In contrast, our LLM-based approach produces one unique solution per iteration without population-based search. Although the maximum number of generations for both GA and OmeGA was set to 2,000 (refer to Appendix 2), the convergence comparison in Figure 4 focuses on the first 10,000 unique solutions to ensure clarity of visualization. The shaded bands around each curve in Figure 4 indicate one standard deviation across multiple independent runs.
Across all four DSM cases, our methods consistently demonstrate superior convergence speed relative to all GA variants and remain highly competitive with the purpose-specific OmeGA algorithm. The LLM-based methods reach substantially lower objective values with far fewer explored sequences, highlighting their efficiency under constrained iteration budgets. Compared to OmeGA, our LLM-based methods often identify high-quality solutions earlier and continue refining them through iterative improvement. This observation supports the hypothesis that LLMs can perform efficient in-context learning, rapidly adapting and optimizing based on previously explored sequences. The only exception occurs in the Automobile Brake System DSM, where OmeGA achieves the optimal solution from the very beginning due to the relatively simple structure of the DSM’s strongly connected components. In this case, the block decomposition provided by the SCCs drastically reduces the search space, enabling the genetic algorithm with local search to immediately converge to the global optimum.
@@
Moreover, the experimental results further highlight the value of contextual domain knowledge in improving the convergence behavior. In Figures 4A (UCAV), 4C (Heat Exchanger), and 4D (Brake System), the version of our method that incorporates domain knowledge (red line) consistently outperforms the variant without such knowledge (gray line), both in terms of convergence speed and final solution quality. This indicates that LLMs are capable of effectively combining semantic reasoning with structural optimization to solve CO problems more intelligently. A noteworthy exception appears in Figure 4B (Microfilm Cartridge), where both variants of the LLM-based method reach the optimal solution in the first step. As a result, the two lines overlap entirely, reflecting identical performance in this case. The findings indicate that for some DSMs with relatively less complexity, a small number of attempts may be sufficient for the LLM to infer optimal solutions.
@@ 5.4 Implications for Engineering Design
Figure 5: Optimization trajectory for the Heat Exchanger DSM case using the proposed method
@@
In sum, the proposed method can directly contribute to design planning and system-level coordination in engineering design practice. For example, in activity-based DSMs, it supports early-phase decision-making by suggesting task sequences that mitigate rework and streamline communication. In parameter-based DSMs, it reduces technical complexity by identifying sequencing patterns that minimize feedback loops. By leveraging both system structure information and domain knowledge, our method enables the LLM to reason in alignment with engineering contexts, rather than relying solely on statistical correlations.
While this study focuses on sequencing problems within DSMs, the proposed approach can be generalized to a broader class of CO problems [63, 64, 65, 66, 67] in engineering and more fields. A more generalized framework is illustrated in Figure 6. During the engineering design process, systems or tasks can often be formulated as iterative optimization problems over formal representations, where multimodal information (such as task descriptions, parameter interactions, or architectural constraints encoded in text, diagrams, or code format) is encoded and integrated. The framework captures this pattern by transforming such information into a combined input information and integrating them with objective evaluation in a closed-loop reasoning cycle, enabling generation-based optimization using AI foundation models.
Figure 6: Foundation Model-based Framework for Solving General CO Problems
@@ 6 Limitations and Future Work
While the proposed method demonstrates strong performance, several limitations remain. First, DSMs in practice are highly diverse, not only in terms of size and structural complexity, but also in their representation. Beyond binary DSMs, weighted DSMs, multi-domain DSMs, and other enriched forms are common, where edge strength, type, or cross-domain dependencies play a critical role [1]. Future work should evaluate the LLM-based methodology on a broader spectrum of DSM types to better reflect real engineering settings. Second, the scalability of the method to large-scale DSMs remains an open question. As system size grows, both computational performance and solution quality may be affected. While the proposed framework benefits from the flexible reasoning capability of LLMs, further work is needed to systematically assess its efficiency and convergence behavior on large-scale networks. Third, our current evaluation uses the minimization of feedback loops as the sequencing objective. However, prior studies have proposed a wide variety of objectives, including reducing feedback length, to balancing concurrency, to minimizing project cost and duration, and to enhancing modularity [51, 49, 48, 2]. Future work should extend testing to these alternative objectives and different classes of combinatorial optimization problems beyond sequencing, such as modularization and change propagation analysis [63, 64, 65, 66, 67].
Another important direction for future research lies in improving the interpretability of the optimization process. In this paper, we presented a preliminary case illustrating how the LLM refines the DSM through a sequence of iterative updates. However, the underlying decision-making and reasoning mechanisms of LLMs still remain relatively opaque. Researchers may explore more effective techniques for LLM interpretability, such as integrating complementary models, enabling self-explanation, and visualizing decision pathways [68]. Additionally, the current framework primarily relies on textual representations of domain knowledge and system structures. While effective, this may limit its ability to fully leverage non-textual modalities such as diagrams, tables, or images commonly used in engineering practice [69, 70]. In future work, we will explore Multimodal Large Language Models (MLLMs) to incorporate visual or tabular information [71], enriching the reasoning input and expanding the framework’s applicability to more complex engineering workflows.
In addition, DSM methods have already been incorporated into commercial tools such as Loomeo (https://loomeo.com/) and Lattix (https://www.lattix.com/), which provide standardized DSM representations, built-in objectives, and user-friendly interfaces. Future studies may explore the comparison between our proposed method and these commercial tools. Moreover, building on this foundation, LLM-based optimization could be integrated into such platforms to strengthen their adaptability and intelligence. A potential implementation would involve using the existing DSM representation layer and evaluation functions as the infrastructure, while adding an LLM-based optimization module that generates and iteratively refines candidate sequences through prompt templates (refer to Appendix 1). The outputs from the LLM could then be automatically validated against the software’s scoring functions, and retrieval-augmented generation could be employed to incorporate external knowledge sources (e.g., design guidelines, prior project documentation) to enrich the reasoning process.
@@ 7 Conclusion
In this paper, we proposed a knowledge-informed optimization approach based on LLMs for solving DSM sequencing, a representative combinatorial optimization problem in engineering design. By integrating structural network information with contextual domain knowledge, the approach exploits the reasoning capabilities of LLMs to iteratively generate and refine candidate solutions. Experiments on various DSM cases show that the proposed method outperforms both stochastic and deterministic baselines in terms of convergence speed and solution quality. The consistent improvements observed when domain knowledge is incorporated highlight the benefit of combining semantic insight with mathematical structure. Overall, this study demonstrates that LLMs provide a promising foundation for a new paradigm in engineering design optimization and decision support, particularly in complex and knowledge-intensive engineering contexts.
@@ Appendix 1. Full prompts
Prompt for the proposed approach (without knowledge):
@@
Appendix 2. Parameter Settings for Stochastic Baseline Methods
In all three variants of the GA used for DSM sequencing tasks, we employed the following shared settings: the number of generations was set to 2,000, with a selection mechanism using tournament selection and mutation using shuffled indexes. Algorithms were implemented using the DEAP library with standard configurations for initial population generation. To ensure a fair comparison, all three GA variants and the OmeGA algorithm were initialized using the same seed solution as the proposed LLM-based methods. The different configurations for each GA setting are shown in Table 4 below. The detailed description of each parameter refers to [72].
Table 4: Parameter settings for three GA variants
@@ Appendix 2. Parameter Settings for Stochastic Baseline Methods
For the experimental setting of the Ordered Messy Genetic Algorithm (OmeGA) [36], the DSM was first decomposed into strongly connected components using depth-first search algorithm [73], followed by topological sorting to determine inter-block ordering. Within each block, the GA was run with the same base settings as above but integrated with a 2-opt local search for further refinement. The population size per block was set to 50, the number of generations to 2,000, the tournament size to 4, crossover probability to 0.7, and mutation probability to 0.2. The local search was applied to the best individual in each generation with a maximum of 20 iterations. These parameter settings follow the literature with minor modifications [36].